Solve by factoring.
x^2+8x-20=0

Answer by edjones(8007)

x^2+8x-20=0
The last sign tells us the signs are going to be different.
The middle sign is the sign of the larger number.
Possible factors of 20:
20 -1=+19
10 -2=+8
5 -4=+1
(x+10)(x-2)
Ed